OpenStudy (anonymous):

ca-6a+cb-6b factoring polynomials with more then three terms?? I'm looking for some sort of step by step instructions on how go about solving these !!!

OpenStudy (anonymous):

First thing you do is group like terms... so you have ca-6a = a(c-6) and cb-6b=b(c-6) this leaves you with... a(c-6)+b(c-6) Just like you did before you can factor out the term (c-6) giving you... (c-6)(a+b)